OpenStudy (lgbasallote):

Solve without using calculator. Use two decimal places. \[\sqrt{200} - \sqrt{32}\]

OpenStudy (lgbasallote):

please just show how

OpenStudy (shubhamsrg):

sqrt(200) = 10sqrt(2) and sqrt(32) = 4sqrt(2) sqrt(2) = 1.414 approx.. easy now?

Parth (parthkohli):

Okay, step by step, factor both the terms.

OpenStudy (lgbasallote):

how did you get the approximation?

OpenStudy (auctoratrox):

sqrt(2x100) - sqrt (2x16) = 10sqrt2 - 4sqrt2 = 6sqrt2

OpenStudy (lgbasallote):

well yeah i get \[6\sqrt2\] but how to solve that?

Parth (parthkohli):

Now factor \(\sqrt 32\)

OpenStudy (lgbasallote):

i just want to know the value of \(6\sqrt 2\) without calculator use -_-

OpenStudy (auctoratrox):

I always remember sqrt2 = 1.4something, not sure about this one

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow \sqrt{32} = 4\sqrt{2} }\)

OpenStudy (lgbasallote):

lol not good enough

OpenStudy (lgbasallote):

how could i verify it is 1.4

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 6\sqrt{2} - 4\sqrt{2} = 2\sqrt{2} }\)

OpenStudy (auctoratrox):

Are you taking calculus? you could use a linear approximation to get it if you know how

OpenStudy (lgbasallote):

yeah...that...i need to know the formula for linear approx

Parth (parthkohli):

You can find square roots using division method.

OpenStudy (lgbasallote):

and how to use it

Parth (parthkohli):

See the long division method.

OpenStudy (lgbasallote):

idk what you're talking about @ParthKohli o.O

OpenStudy (auctoratrox):

I definitely didn't learn that method, and guess and check sucks. not sure what to tell you

Parth (parthkohli):

Man!

Parth (parthkohli):

The division method is what I am talking about.

OpenStudy (lgbasallote):

nevermind then haha

OpenStudy (lgbasallote):

seems no one knows linear approximation

OpenStudy (ajprincess):

see between which square numbers 2 lies. it lies between 1 and 4 see to which number 2 is closer. it is closer to 1 than 4. nw sqrt of 1 is 1 sqrt of 4 is 2 so sqrt of 2 must be closer to 1 square 1.1, 1.2, 1.3, 1.4 and see whose square is approximately 2 . then that is the sqrt of 2. hope that helps.

OpenStudy (shubhamsrg):

well to find x^1/2 ,,you can use calculus,,let f(x) = x^1/2 take a no. close to 2 whose sqrt is known to you let x=1.96 (as sqrt(1.96) =1.4) { you can also take 2.25 as sqrt(2.25) = 1.5} dx = 2 - 1.96 = 0.04 now according to lebinitz rules,, dy = f(x+dx) - f(x) thus dy = 2^1/2 - 1.96^1/2 also,,dy = (f'(x))dx thus dy = 1/(2x^1/2) (dx) = 1/(2*1.4) ( 0.04) = 0.04/2.8 = 1/70 substitute it in 1st eqn thus 2^1/2 = 1/70 + 1.4 = 1.414 .... hope it was helpful..

OpenStudy (shubhamsrg):

note that above approx could be made as x>>dx ..